# Quantitative surgery and total mean curvature

### v1

@@ 4. Surgery bending in Euclidean space
$\lambda_{2}=\dots=\lambda_{n-k}=-4\,\lambda_{1}.$
on the graph of $f_{\varrho}$ on $\varrho^{4},\sqrt{\varrho}]$. Therefore, we get $\lambda_{1}=-(16(n-k)-15)^{-1/2}|A|$ and thus
$H(\overline{\Sigma}_{\varrho})$ $=\lambda_{1}-4(n-k-1)\lambda_{1}=\frac{4(n-k)-5}{\sqrt{16(n-k)-15}}|A|$
$>\frac{1}{2}|A|$ $\text{if }n-k\geq 2$
$\mathrm{scal}(\overline{\Sigma}_{\varrho})$ $=16(n-k-1)(n-k-2)\lambda_{1}^{2}-8(n-k-1)\lambda_{1}^{2}$
$=8(n-k-1)(2(n-k)-5)\lambda_{1}^{2}$
$=\frac{8\left((n-k)-1\right)\left(2(n-k)-5\right)}{16(n-k)-15}|A|^{2}$
$>\frac{1}{3}|A|^{2}$ $\text{if }n-k\geq 3.$
On the other hand, the cylinder of radius $\varrho^{4}$ has principal curvatures: $\lambda_{1}=0$ and $\lambda_{2}=\dots=\lambda_{n-k}=\varrho^{-4}\not=0$ and therefore
(4.2) $\begin{aligned} H(\overline{\Sigma}_{\varrho})&=\sqrt{n-k-1}\,|A|>\frac{1}{2}|A|&&\text{if }n-k\geq 2\\ \mathrm{scal}(\overline{\Sigma}_{\varrho})&=(n-k-2)|A|^{2}>\frac{1}{3}|A|^{2}&&\text{if }n-k\geq 3.\end{aligned}$
Next, we show (iii). On the cylindrical part, we have $|A|=\sqrt{n-k-1}\varrho^{-4}$, so the desired inequality holds by 4.2. On the “graphical part”, we observe that both $\xi_{\varrho}^{\prime}$ and the function $x\mapsto\frac{-x}{\sqrt{1+x^{2}}}$ are monotone decreasing and thus, for $r\leq\varrho^{2}$:
@@
$H={}$ $-\frac{1}{4}\lambda_{2}+(n-k-1)\lambda_{2}=(n-k-\tfrac{5}{4})\lambda_{2}$
$\geq{}$ $(n-k-\tfrac{5}{4})\varrho^{-3/2}$
$\mathrm{scal}={}$ $(n-k-1)(n-k-2)\lambda_{2}^{2}-\frac{1}{2}(n-k-1)\lambda_{2}^{2}$
$\geq{}$ $(n-k-1)(n-k-\tfrac{5}{2})\varrho^{-3}$
and if $n-k\geq 3$, then $H>\varrho^{-3/2}$ and $\mathrm{scal}\geq\varrho^{-3}$, which proves (iii).
@@
$0\geq f_{\varrho}^{\prime}(r)\geq\xi_{\varrho}^{\prime}(\varrho^{2})=-\frac{\varrho}{\sqrt{\varrho-\varrho^{2}}}\geq-2\sqrt{\varrho}.$
@@
$|\overline{\nu}_{\varrho}^{\intercal}|=\frac{|\nabla f_{\varrho}|}{\sqrt{1+|\nabla f_{\varrho}|^{2}}}\leq|\nabla f_{\varrho}|=|f_{\varrho}^{\prime}|\leq 2\sqrt{\varrho}.$
@@
$\xi_{\varrho}^{\prime}(r)=-\frac{\varrho}{\sqrt{\sqrt{r}-\varrho^{2}}},\qquad 4\xi_{\varrho}^{\prime\prime}(r)=-\frac{1}{r}\xi_{\varrho}^{\prime}(r)\left(1+\xi_{\varrho}^{\prime}(r)^{2}\right).$
@@
$\lambda_{2}=\dots=\lambda_{n-k}={}$ $-\frac{f_{\varrho}^{\prime}}{r\sqrt{1+f_{\varrho}^{\prime 2}}}{\leq}\frac{3\sqrt{2}\,\varrho^{7/8}}{\sqrt{\varrho}}=3\sqrt{2}\,\varrho^{3/8}$
@@
(4.3) $|A(\overline{\Sigma}_{\varrho})|(r)=\sqrt{\lambda_{1}^{2}+(n-k-1)\lambda_{2}^{2}}\leq 3\sqrt{2(n-k+24)}\,\varrho^{3/8}.$
The scalar curvature estimate can be derived from the Gauß-equation
@@
which together with (4.3) yields the desired inequality and hence concludes the proof of (v). Lastly, we point out that the analysis still holds after smoothing out the $\overline{\Sigma}_{\varrho}$ at the gluing region $t=a_{\varrho}$, see for instance [3]. ∎
@@
Lemma 4.4.
For every sufficiently small $R$ and $\varrho_{1}\in(0,R^{2}/4]$, there exists a family of diffeomorphisms $\Phi_{\varrho_{1},\varrho}:\overline{\Sigma}_{\varrho_{1}}\to\overline{\Sigma}_{\varrho}$, $\varrho\in(0,\varrho_{1}]$, such that $\Phi_{\varrho_{1},\varrho_{1}}=\operatorname{\mathrm{id}}_{\overline{\Sigma}_{\varrho_{1}}}$ and for the metrics $g_{\varrho,\delta}$ on $\overline{\Sigma}_{\varrho}$ induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$, the assignment
@@
Before we can prove Lemma 4.4, we need some preparation. Since the Euclidean handle $\overline{\Sigma}_{\varrho}\subseteq B_{1}^{n-k}(0)\times[0,1]$ is rotationally symmetric, it can be obtained by rotating a curve $\gamma_{\varrho}$. Hence, we may write the metric $g_{\overline{\Sigma}_{\varrho}}$ on $\overline{\Sigma}_{\varrho}$ induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$ in warped product form
$g_{\varrho,\delta}=\mathrm{d}\tau^{2}+r_{\varrho}(\tau)^{2}g_{S^{n-k-1}},$
where $\tau$ is the arc-length along the generating curve $\gamma_{\varrho}$ and $r_{\varrho}(\tau)$ is the radius-function. For a fixed $R\in(0,\tfrac{2}{\sqrt{10}}]$ and all $\varrho<R^{2}/4$, we denote by $d_{\varrho}$ the length of the curve $\gamma_{\varrho}$ until radius $R$. Since $\Sigma_{\varrho}$ ends in a straight cylinder, the radial function satisfies $r_{\varrho}|_{[d_{\varrho}-1+a_{\varrho},d_{\varrho}]}=\varrho^{4}$. We have the following estimate on $d_{\varrho}$ and its $\varrho$-derivative.
@@
There exists a sufficiently small $R>0$ depending only on $n$, such that for every $\varrho<R^{2}/4$, we have
@@
Recall the function $f_{\varrho}(r)\coloneqq\eta_{\varrho}(r)\bigl(\xi_{\varrho}(r)+a_{\varrho}\bigr)$. The length $d_{\varrho}$ of $\gamma_{\varrho}$ is given
$d_{\varrho}=\int_{\varrho^{4}}^{4\varrho^{4}}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr+\int_{4\varrho^{4}}^{R}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr+(1-a_{\varrho})$
@@
We estimate $A_{0}$, $A_{1}$ and $A_{2}$ and $\partial_{\varrho}a_{\varrho}$ individually, starting with $A_{0}$. We note that an anti-derivative of $\sqrt{1+f^{\prime}_{\varrho}(r)^{2}}$ is given by
@@
$\leq$ $\Big(|\eta_{\varrho}^{\prime}||\xi_{\varrho}+a_{\varrho}|+|\eta_{\varrho}||\xi_{\varrho}^{\prime}|\Big)\cdot\Big(|\partial_{\varrho}\eta_{\varrho}^{\prime}||\xi_{\varrho}+a_{\varrho}|+|\partial_{\varrho}\eta_{\varrho}||\xi_{\varrho}^{\prime}|+|\eta_{\varrho}^{\prime}||\partial_{\varrho}\xi_{\varrho}+\partial_{\varrho}a_{\varrho}|+|\eta_{\varrho}||\partial_{\varrho}\xi_{\varrho}^{\prime}|\Big)$
$\lesssim$ $\Big(\varrho^{-1/2}\cdot\varrho^{5/4}+1\cdot\varrho^{3/4}\Big)\cdot\Big(\varrho^{-2}\cdot\varrho^{5/4}+\varrho^{-3/2}\cdot\varrho^{3/4}+\varrho^{-1/2}\cdot(1+1)+1\cdot\varrho^{-1/8}\Big)$
$\lesssim$ $\varrho^{3/4}\cdot\varrho^{-3/4}.$
@@
Since the curve $\gamma_{\varrho}$ is shorter than the one consisting of staight parallel to the coordinate-axes and longer than the straight line connecting its end points, we get
@@
Proof of Lemma 4.4.
Fix the bump function $\alpha(x)=\eta(4x)\eta(-4x)$, where $\eta$ is the cut-off function introduced earlier satisfying $2\leq\eta^{\prime}\leq 0$ and $|\eta^{\prime\prime}|<10$ and we note that $\operatorname{supp}\alpha\subseteq[-\frac{1}{4},\frac{1}{4}]$. Next, we define a smooth family or reparameterizations $\varphi_{\varrho_{1},\varrho}:[0,d_{\varrho_{1}}]\to[0,d_{\varrho}]$ for $\varrho\in(0,\varrho_{1}]$ by
$\varphi_{\varrho_{1},\varrho}(\tau)=\tau-q_{\varrho_{1},\varrho}\int_{0}^{\tau}\alpha\left(-d_{\varrho}+\tfrac{1}{4}+s\right)ds$
where $q_{\varrho_{1},\varrho}=(d_{\varrho_{1}}-d_{\varrho})\left(\int_{\mathbb{R}}\alpha(s)ds\right)^{-1}$. We now define a diffeomorphism
@@
$\overline{\Phi}_{\varrho_{1},\varrho}^{*}g_{\varrho,\delta}(\tau,x)=(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau))^{2}\mathrm{d}\tau^{2}+r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))^{2}g_{S^{n-k-1}}.$
@@
$\partial_{\tau}\varphi(\tau)=1-(d_{\varrho_{1}}-d_{\varrho})\alpha\left(-d_{\varrho}+\tfrac{1}{4}+\tau\right)$
@@
$\partial_{\varrho}\partial_{\tau}\varphi(\tau)=\alpha(-d_{\varrho}+\tfrac{1}{4}+\tau)\partial_{\varrho}d_{\varrho}-(d_{\varrho_{1}}-d_{\varrho})\alpha^{\prime}(-d_{\varrho}+\frac{1}{4}+\tau)\partial_{\varrho}d_{\varrho}>-\frac{1}{2}\partial_{\tau}\varphi(\tau)$
where we used that $(d_{\varrho_{1}}-d_{\varrho})\ll 1$ and that $\alpha^{\prime}$ is bounded. Consequently, we obtain for $\varrho\in[\varrho_{0},\varrho_{1}]$

### v2

@@ 5. Surgery bending in Euclidean space
$\lambda_{2}=\dots=\lambda_{k}=-4\,\lambda_{1}.$
On the graph of $f_{\varrho}$ on $\varrho^{4},\sqrt{\varrho}]$. Therefore, we get $\lambda_{1}=-(16k-15)^{-1/2}|A|$ and thus
$H(\overline{\Sigma}_{\varrho})$ $=\lambda_{1}-4(k-1)\lambda_{1}=\frac{4k-5}{\sqrt{16k-15}}|A|$
$>\frac{1}{2}|A|$ $\text{if }k\geq 2$
$\mathrm{scal}(\overline{\Sigma}_{\varrho})$ $=16(k-1)(k-2)\lambda_{1}^{2}-8(k-1)\lambda_{1}^{2}$
$=8(k-1)(2k-5)\lambda_{1}^{2}$
$=\frac{8\left(k-1\right)\left(2k-5\right)}{16k-15}|A|^{2}$
$>\frac{1}{3}|A|^{2}$ $\text{if }k\geq 3.$
On the other hand, the cylinder of radius $\varrho^{4}$ has principal curvatures: $\lambda_{1}=0$ and $\lambda_{2}=\dots=\lambda_{k}=\varrho^{-4}\not=0$ and therefore
(5.2) $\begin{aligned} H(\overline{\Sigma}_{\varrho})&=\sqrt{k-1}\,|A|>\frac{1}{2}|A|&&\text{if }k\geq 2\\ \mathrm{scal}(\overline{\Sigma}_{\varrho})&=(k-2)|A|^{2}>\frac{1}{3}|A|^{2}&&\text{if }k\geq 3.\end{aligned}$
Next, we show (iii). On the cylindrical part, we have $|A|=\sqrt{k-1}\varrho^{-4}$, so the desired inequality holds by (5.2). On the “graphical part”, we observe that both $\xi_{\varrho}^{\prime}$ and the function $x\mapsto\frac{-x}{\sqrt{1+x^{2}}}$ are monotone decreasing and thus, for $r\leq\varrho^{2}$:
@@
$H={}$ $-\frac{1}{4}\lambda_{2}+(k-1)\lambda_{2}=(k-\tfrac{5}{4})\lambda_{2}$
$\geq{}$ $(k-\tfrac{5}{4})\varrho^{-3/2}$
$\mathrm{scal}={}$ $(k-1)(k-2)\lambda_{2}^{2}-\frac{1}{2}(k-1)\lambda_{2}^{2}$
$\geq{}$ $(k-1)(k-\tfrac{5}{2})\varrho^{-3}$
and if $k\geq 3$, then $H>\varrho^{-3/2}$ and $\mathrm{scal}\geq\varrho^{-3}$, which proves (iii).
@@
$0\geq f_{\varrho}^{\prime}(r)\geq\xi_{\varrho}^{\prime}(\varrho^{2})=-\frac{\varrho}{\sqrt{\varrho-\varrho^{2}}}\geq-2\sqrt{\varrho},$
@@
$|\overline{\nu}_{\varrho}^{\intercal}|=\frac{|\nabla f_{\varrho}|}{\sqrt{1+|\nabla f_{\varrho}|^{2}}}\leq|\nabla f_{\varrho}|=|f_{\varrho}^{\prime}|\leq 2\sqrt{\varrho},$
@@
$\xi_{\varrho}^{\prime}(r)=-\frac{\varrho}{\sqrt{\sqrt{r}-\varrho^{2}}},\qquad 4\xi_{\varrho}^{\prime\prime}(r)=-\frac{1}{r}\xi_{\varrho}^{\prime}(r)\left(1+\xi_{\varrho}^{\prime}(r)^{2}\right),$
@@
$\lambda_{2}=\dots=\lambda_{k}={}$ $-\frac{f_{\varrho}^{\prime}}{r\sqrt{1+f_{\varrho}^{\prime 2}}}{\leq}\frac{3\sqrt{2}\,\varrho^{7/8}}{\sqrt{\varrho}}=3\sqrt{2}\,\varrho^{3/8}$
@@
(5.3) $|A(\overline{\Sigma}_{\varrho})|(r)=\sqrt{\lambda_{1}^{2}+(k-1)\lambda_{2}^{2}}\leq 3\sqrt{2(k+24)}\,\varrho^{3/8}.$
The scalar curvature estimate can be derived from the Gauss equation
@@
which together with (5.3) yields the desired inequality and hence concludes the proof of (v). Lastly, we point out that the analysis still holds after smoothing out the $\overline{\Sigma}_{\varrho}$ at the gluing region $t=a_{\varrho}$, see for instance [BH22]. ∎
@@
Proposition 5.4.
For every sufficiently small $R\in(0,\tfrac{2}{\sqrt{10}}]$ and $\varrho_{1}\in(0,R^{2}/4]$, there exists a smooth path of diffeomorphisms $\Phi_{\varrho_{1},\varrho}:\overline{\Sigma}_{\varrho_{1}}\to\overline{\Sigma}_{\varrho}$, $\varrho\in(0,\varrho_{1}]$, such that $\Phi_{\varrho_{1},\varrho_{1}}=\operatorname{\mathrm{id}}_{\overline{\Sigma}_{\varrho_{1}}}$ and for the metrics $g_{\varrho,\delta}$ on $\overline{\Sigma}_{\varrho}$ induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$, the assignment
@@
Before we can prove Proposition 5.4, we need some preparation. Since the Euclidean handle $\overline{\Sigma}_{\varrho}\subseteq B_{1}^{k}(0)\times[0,1]$ is rotationally symmetric, it can be obtained by rotating a curve $\overline{\gamma}_{\varrho}$. Hence, we may write the metric $g_{\overline{\Sigma}_{\varrho}}$ on $\overline{\Sigma}_{\varrho}$ induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$ in warped product form
(5.4) $g_{\varrho,\delta}=d\tau^{2}+r_{\varrho}(\tau)^{2}g_{S^{k-1}},$
where $\tau$ is the arc-length along the generating curve $\overline{\gamma}_{\varrho}$ and $r_{\varrho}(\tau)$ is the radius-function. The parametrization is such that $r_{\varrho}$ is monotonically non-increasing. For $R\in(0,\tfrac{2}{\sqrt{10}}]$ and all $\varrho<R^{2}/4$, we denote by $d_{\varrho}$ the length of the curve $\overline{\gamma}_{\varrho}$ until radius $R$. Since $\Sigma_{\varrho}$ ends in a straight cylinder, the radial function satisfies $r_{\varrho}|_{[d_{\varrho}-1+a_{\varrho},d_{\varrho}]}=\varrho^{4}$. We have the following estimate on $d_{\varrho}$ and its $\varrho$-derivative.
@@
There exists $R\in(0,\tfrac{2}{\sqrt{10}}]$, depending only on $n$, such that for every $\varrho<R^{2}/4$, we have
@@
Recall the function $f_{\varrho}(r)\coloneqq\eta_{\varrho}(r)\bigl(\xi_{\varrho}(r)+a_{\varrho}\bigr)$. The length $d_{\varrho}$ of $\overline{\gamma}_{\varrho}$ is given
$d_{\varrho}=\int_{\varrho^{4}}^{4\varrho^{4}}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr+\int_{4\varrho^{4}}^{R}\sqrt{1+f_{\varrho}^{\prime}(r)^{2}}dr+(1-a_{\varrho}),$
@@
We estimate $A_{0}$, $A_{1}$ and $A_{2}$ and $\partial_{\varrho}a_{\varrho}$ individually, starting with $A_{0}$. We note that an antiderivative of $\sqrt{1+f^{\prime}_{\varrho}(r)^{2}}$ is given by
@@
$\quad\leq\Big(|\eta_{\varrho}^{\prime}||\xi_{\varrho}+a_{\varrho}|+|\eta_{\varrho}||\xi_{\varrho}^{\prime}|\Big)$
$\qquad\qquad\cdot\Big(|\partial_{\varrho}\eta_{\varrho}^{\prime}||\xi_{\varrho}+a_{\varrho}|+|\partial_{\varrho}\eta_{\varrho}||\xi_{\varrho}^{\prime}|+|\eta_{\varrho}^{\prime}||\partial_{\varrho}\xi_{\varrho}+\partial_{\varrho}a_{\varrho}|+|\eta_{\varrho}||\partial_{\varrho}\xi_{\varrho}^{\prime}|\Big)$
$\quad\lesssim\Big(\varrho^{-1/2}\cdot\varrho^{5/4}+1\cdot\varrho^{3/4}\Big)$
$\qquad\qquad\cdot\Big(\varrho^{-2}\cdot\varrho^{5/4}+\varrho^{-3/2}\cdot\varrho^{3/4}+\varrho^{-1/2}\cdot(1+1)+1\cdot\varrho^{-1/8}\Big)$
$\quad\lesssim\varrho^{3/4}\cdot\varrho^{-3/4},$
@@
Since the curve $\overline{\gamma}_{\varrho}$ is shorter than the one consisting of straight ones parallel to the coordinate-axes and longer than the straight line connecting its end points, we get
@@
Proof of Proposition 5.4.
Fix the bump function $\beta(x)=\eta(4x)\eta(-4x)$, where $\eta$ is the cut-off function introduced earlier satisfying $2\leq\eta^{\prime}\leq 0$ and $|\eta^{\prime\prime}|<10$, and we note that $\operatorname{supp}\beta\subseteq[-\frac{1}{4},\frac{1}{4}]$. Next, we define a smooth family or reparameterizations $\varphi_{\varrho_{1},\varrho}:[0,d_{\varrho_{1}}]\to[0,d_{\varrho}]$ for $\varrho\in(0,\varrho_{1}]$ by
(5.5) $\varphi_{\varrho_{1},\varrho}(\tau)=\tau-q_{\varrho_{1},\varrho}\int_{0}^{\tau}\beta\left(s-\frac{1}{2}\right)ds$
where $q_{\varrho_{1},\varrho}=(d_{\varrho_{1}}-d_{\varrho})\left(\int_{\mathbb{R}}\beta(s)ds\right)^{-1}$. We now define a diffeomorphism
@@
$\overline{\Phi}_{\varrho_{1},\varrho}^{*}g_{\varrho,\delta}(\tau,x)=(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau))^{2}\mathrm{d}\tau^{2}+r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))^{2}g_{S^{k-1}}.$
@@
$\partial_{\tau}\varphi(\tau)=1-(d_{\varrho_{1}}-d_{\varrho})\beta\left(\tau-\frac{1}{2}\right)$
@@
$\partial_{\varrho}\partial_{\tau}\varphi(\tau)=\beta\left(\tau-\frac{1}{2}\right)\partial_{\varrho}d_{\varrho}-(d_{\varrho_{1}}-d_{\varrho})\beta^{\prime}(\tau-\frac{1}{2})>-\frac{1}{2}\partial_{\tau}\varphi(\tau)$
where we used that $(d_{\varrho_{1}}-d_{\varrho})\ll 1$ and that $\beta^{\prime}$ is bounded. Consequently, we obtain for $\varrho\in[\varrho_{0},\varrho_{1}]$
@@
To conclude the construction of $\overline{\Sigma}$, we will now smoothen the $C^{1}$-singularity of $\overline{\Sigma}_{\varrho}$ at $\{t=a_{\varrho}\}$ while preserving Proposition 5.3 and Lemma 5.5, hence also Proposition 5.4. Fix some $R\in(0,\tfrac{2}{\sqrt{10}}]$ so that Lemma 5.5 holds for this $R$. Consider the smooth manifold
$V=\{(\varrho,t)\mid\varrho\in(0,R^{2}/4),t\in(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}\subset(0,R^{2}/4)\times(0,1]$
with non-compact boundary $\partial V=\{(\varrho,a_{\varrho})\mid\varrho\in(0,R^{2}/4)\}\subset V$.
Put $f_{0}\colon V\to\mathbb{R}$, $f_{0}(\varrho,t)=\xi_{\rho}^{-1}(t-a_{\varrho})$. Note that for fixed $\varrho$, the function $f_{0}(\varrho,-)\colon(f_{\varrho}(\varrho^{2}),a_{\varrho}]\to\mathbb{R}$ is the inverse of the function $\varrho^{4},\varrho^{2})\to\mathbb{R}$, $r\mapsto f_{\varrho}(r)$.
Let $\omega\colon V\to\mathbb{R}_{>0}$ be a smooth function. For each $\varrho\in(0,R^{2}/4)$, we obtain the surface of revolution
$S(\omega,\varrho):=\{r=\omega(t)\mid t\in(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}\subset\mathbb{R}^{n}\times(f_{\varrho}(\varrho^{2}),a_{\varrho}].$
Let $J^{2}V\to V$ be the bundle of $2$-jets of strictly positive real valued smooth functions on $V$. We consider the open partial differential relation $\mathfrak{R}\subset J^{2}V$ determined by the following pointwise conditions for $\omega$, working with the numbers $|\partial_{\varrho}d_{\varrho}|<\varrho^{1/4}$, from Lemma 5.5:
  1. (i)
    $H(S(\omega,\varrho))>\frac{1}{2}|A(S(\omega,\varrho))|$ if $k\geq 2$,
  2. (ii)
    $\mathrm{scal}(S(\omega,\varrho))>\frac{1}{3}|A(S(\omega,\varrho))|^{2}$ if $k\geq 3$,
  3. (iii)
    $\min\left\{H(S(\omega,\varrho))^{2},\mathrm{scal}(S(\omega,\varrho))\right\}>\tfrac{1}{2}\varrho^{-3}$
  4. (iv)
    $\left|\partial_{\varrho}\left(\sqrt{1+\omega^{\prime}(\varrho,-)}-\sqrt{1+f^{\prime}_{0}(\varrho,-)}\right)\right|<\varrho^{1/4}-|\partial_{\varrho}d_{\varrho}|$.
By Proposition 5.3, $f_{0}$ solves $\mathfrak{R}$ over $V$. Consider the smooth path $F\colon[0,1]\to C^{\infty}(V,\mathbb{R}_{>0})$,
$F(\tau)(\varrho,t):=(1-\tau)f_{0}(\varrho,t)+\tau\varrho^{4}.$
We have $j^{1}F(\tau)|_{\partial V}=j^{1}f_{0}|_{\partial V}$ for all $t$. Carrying out the computation in the proof of part (ii) and (iii) of Proposition 5.3 with $\lambda_{2}=\ldots=\lambda_{k}=\varrho^{4}$ and $\lambda_{1}\in[-\tfrac{1}{4}\varrho^{4},0]$, one sees that $F$ solves $\mathfrak{R}$ along $\partial V$ for all $\tau$ (for $\omega=F(\tau)$, the left-hand side of (iv) is equal to 0 along $\partial V$). By continuity, there exists an open neighborhood $\partial V\subset U\subset V$ such that $F$ solves $\mathfrak{R}$ over $U$ for all $t$.
By the local flexibility lemma [BH22, Theorem 1.2] applied to the closed subset $\partial V\subset V$, there exists an open neighborhood $\partial V\subset U_{0}\subset U$ satisfying $\overline{U}_{0}\subset V$ and a smooth path $f\colon[0,1]\to C^{\infty}(V,\mathbb{R}_{>0})$ such that $f(0)=f_{0}$, $f(\tau)=F(\tau)$ on $U_{0}$, $f(\tau)=f_{0}$ on $V\setminus U$ and $f(\tau)$ solves $\mathfrak{R}$ over $V$ for all $\tau$.
For $\varrho\in(0,R^{2}/4)$, define
  • the smooth function $\psi_{\varrho}\colon(f_{\varrho}(\varrho^{2}),a_{\varrho}]\to\mathbb{R}_{>0}$, $\psi_{\varrho}(t)=f(1)(\varrho,t)$. It is equal to $\varrho^{4}$ near $a_{\varrho}$ and equal to $f_{0}(\varrho,-)$ near $f_{\varrho}(\varrho^{2})$;
  • the surface of revolution
    $S(\varrho):=\{r=\psi_{\varrho}(t)\mid t\in(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}\subset B_{1}^{k}(0)\times(f_{\varrho}(\varrho^{2}),a_{\varrho}];$
  • the smooth hypersurface
    $\widetilde{\Sigma}_{\varrho}:=\{(r,t)\in\overline{\Sigma}_{\varrho}\mid t\notin(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}\;\bigcup\;S(\varrho)\subset B_{1}^{k}(0)\times[0,1].$
The hypersurface $\widetilde{\Sigma}_{\rho}$ satisfies the properties in Proposition 5.3 by the choice of $\mathfrak{R}$. Note that parts (iv) and (v) of Proposition 5.3 are satisfied since the relevant parts of $\overline{\Sigma}_{\varrho}$ are not affected by the smoothing.
Let $\widetilde{d}_{\rho}$ be defined for $\widetilde{\Sigma}_{\varrho}$ as before Lemma 5.5. Using that $\overline{\Sigma}_{\varrho}=\widetilde{\Sigma}_{\varrho}$ on $\{t\notin(f_{\varrho}(\varrho^{2}),a_{\varrho}]\}$, and that $\psi^{\prime}_{\varrho}(t)=f_{0}^{\prime}(\varrho,-)(t)$ for $t=f_{\varrho}(\varrho^{2})$ and for $t=a_{\varrho}$ (where both sides of the equation are 0), we obtain
$|\partial_{\varrho}\widetilde{d}_{\varrho}|-|\partial_{\varrho}d_{\varrho}|$ $\leq\Big|\partial_{\varrho}(\widetilde{d}_{\varrho}-d_{\varrho})\Big|$
$=\Big|\partial_{\varrho}\int_{f_{\varrho}(\varrho^{2})}^{a_{\varrho}}\left(\sqrt{1+(\psi^{\prime}_{\varrho})^{2}}-\sqrt{1+f_{0}^{\prime}(\varrho,-)^{2}}\right)dt\Big|$
$=\Big|\int_{f_{\varrho}(\varrho^{2})}^{a_{\varrho}}\partial_{\varrho}\left(\sqrt{1+(\psi^{\prime}_{\varrho})^{2}}-\sqrt{1+f_{0}^{\prime}(\varrho,-)^{2}}\right)dt\Big|$
$\leq\big|a_{\varrho}-f_{\varrho}(\varrho^{2})\big|\max_{(f_{\varrho}(\varrho^{2}),a_{\varrho}]}\Big|\partial_{\varrho}\left(\sqrt{1+(\psi^{\prime}_{\varrho})^{2}}-\sqrt{1+f_{0}^{\prime}(\varrho,-)^{2}}\right)\Big|$
$<1\cdot\left(\varrho^{1/4}-|\partial_{\varrho}d_{\varrho}|\right).$
Adding $|\partial_{\varrho}d_{\varrho}|$ to both sides of this inequality proves the first inequality from Lemma 5.5. The inequality $1<\widetilde{d}_{\varrho}$ is clear, while the inequality $\widetilde{d}_{\varrho}<1+R$ is ensured if $\|f(1)-f_{0}\|_{C^{1}(V)}$ is sufficiently small, a condition that can be included in the relation $\mathfrak{R}$.
Finally, we replace $\overline{\Sigma}_{\varrho}$ with the smooth bending handle $\widetilde{\Sigma}_{\varrho}$. By construction, Proposition 5.3 and Lemma 5.5, hold for the new $\overline{\Sigma}_{\varrho}$, which we will use from now on.